OpenStudy (anonymous):

Find the volume of a cone, in terms of pi, with a radius of 3 inches and a height of 6inches. If the radius of the cone is doubled while the height remains constant, what is the volume of the cone? What happens to the volume of a cone if it's radius is doubled?

OpenStudy (anonymous):

V = 1/3 pi*r^2*h r = 3, h = 6 V = 1/3 pi*(3^2)*6 V = 1/3 pi*(9)*(6) V = 1/3 pi*54 V = 18 pi If we double the radius, the volume is four times greater than before. So if our radius goes from 3in to 6in, the volume goes to 72 pi.

OpenStudy (anonymous):

Thank you. I had switched them around and had it wrong.